Find the equation of the line slope1/2 that contains the point(-1,5). please show all work.Ronnorton@adelphia.net
Answer by venugopalramana(3286)

Find the equation of the line slope M=1/2 that contains the point(-1,5)..X1=1...Y1=-5. please show all work
FORMULA IS
Y-Y1=M(X-X1)
Y-(-5)=(1/2)(X-1)
Y+5=(1/2)(X-1)
2Y+10=X-1
2Y=X-11